# Influence of the Galactic Halo on the UHECR Multipoles

### v1

@@ 3 Analysis and results
Figure 4: The skymaps shown above were all produced for $L_{\rm{scat}}\approx 10$ kpc ($B_{\rm{tur}}\approx~{}35~{}\mu$G). Top-left: observed dipole skymap when an extragalactic dipole is injected with the monopole being removed from the skymaps. Top-right: dipole skymap observed when a single beam is the injected source of UHECRs. Bottom-left: observed dipole skymap for the injected three-beam scenario when $L_{\rm{scat}}\approx 10$ kpc. The blue star shows the dipole direction from our simulation and the black circle shows the region provided by PAO for rigidities between 8 to 16 EeV (Abdul Halim et al., 2024). Bottom-right: power spectrum of the observed skymaps for the 3 cases shown above. Note that all maps have a $45^{\circ}$ top-hat smoothing applied (Abreu et al., 2022; Abdul Halim et al., 2024).
To estimate the uncertainties we use the same simulation setup as described in Sec. 2 to push isotropic distributions of particles through the Galaxy but this time without any Galactic magnetic fields or extragalactic dipole. The injected and observed skymaps obtained are used to calculate the dipole and quadrupole amplitudes given by Eq. 3. This noise dipole and quadrupole amplitudes provide a lower limit below which any amplitude calculated is considered to be at the noise level.
The observed skymaps for cases 1, 2, and 3 are shown in the top-right, top-left and bottom-left panel of Fig. 4, respectively. Similar to the injected skymaps the NSIDE = 512 and they have been smoothed by a top-hat function with an angle of $45^{\circ}$ (Abreu et al., 2022; Abdul Halim et al., 2024). For the case when the diffusive sphere in which the UHECRs propagate is comparable or larger than the scattering length of the UHECRs, a large-scale dipolar skymap is obtained. Similar to the injected case we remove the monopole for the single and three-beam observed skymaps. In the three beam case (see bottom-left panel in Fig. 4) the choice of our weights given to the three sources results in observed dipole direction not very far from what is observed by PAO (Abdul Halim et al., 2024). The power spectrum for the three cases is shown in the bottom-right panel of Fig. 4.
In Fig. 5 we examine the observed dipole and quadrupole amplitudes from the three distinct source scenarios and show the following:
@@
    Left panel of Fig. 5: The observed dipole amplitude, $\delta_{0}$, is given by the hollow markers in red, green, and blue for the extragalactic dipole, single and three-beam cases. The dipole is seen to become gradually suppressed as the scattering length, $L_{\rm{scat}}$, is decreased. The solid markers denote the observed dipole amplitude, $\delta_{\rm{all}}$, colour-coded as above, in this case. The dipole amplitude initially increases and later decreases for the single and three beam cases. This behaviour results from the power being unevenly distributed at low values of $\ell$ for large $L_{\rm{scat}}$ values. In appendix A, we plot the power spectrum for the three different scattering length values in Fig. 6 to demonstrate this point. In the extragalactic dipole case, the solid red markers ($\delta_{\rm{all}}$) behave similarly to hollow red markers ($\delta_{0}$), as power at large values of $\ell$ ($>$ 1) is at the noise level. Knowledge of the dipole amplitude alone is insufficient to distinguish between the different source scenarios considered.
@@
    Right panel of Fig. 5. The observed quadrupole amplitude, $\epsilon_{0}$, is given by the hollow markers and $\epsilon_{\rm{all}}$ is given by the solid markers in red, green, and blue for the extragalactic dipole, single and three-beam cases. The quadrupole amplitude obtained for the extragalactic dipole scenario case is very small as we inject only a dipole for these simulations, and hence both the solid and hollow stars in the figure are at the noise level. A non-zero quadrupole amplitude is only obtained for the single and three beam cases.
@@ 4 Discussion
Figure 5: Left-hand panel: dipole amplitude $\delta^{\rm{obs}}$ as a function of scattering length for the three cases: 1) extragalactic dipole shown in red ($\star$), 2) single-beam case shown in green ($\blacksquare$) and 3) three-beam case shown in blue ( ). The solid and hollow markers indicate $\delta_{\rm{all}}$ and $\delta_{0}$, respectively, for all three cases. Right-hand panel: quadrupole amplitude for the three cases as on the left figure with the same colour code, here $\epsilon_{\rm{all}}$ and $\epsilon_{0}$ are shown by solid and hollow markers again. The orange band represents the $1\sigma$ noise of isotropic skymaps with the same statistics (see Sec. 2.2).
@@ 5 Conclusions
[…]
@@ Acknowledgements
VS and AT acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. AvV acknowledges support from Khalifa University’s internal FSU-2022-025 and RIG-2024-047 grants. VS would like to thank Clive Dickinson and the STFC Consolidated Grant (ST/P000649/1) for supporting part of the work carried out at the University of Manchester. The authors would also like to thank James Matthews for useful discussions.
@@ Appendix A Power spectrum for different $L_{\rm{scat}}$ values
We plot the power spectrum from the observed skymaps for three different values of $L_{\rm{scat}}$ in Fig. 6 for the three beam case with $L_{\rm{coh}}\approx$ 150 pc, most of the power sits in the monopole, dipole and quadrupole.

### v2

@@ 3 Analysis and results
[…]
To estimate the uncertainties we use the same simulation setup as described in Sec. 2, but for an isotropic distribution of particles (of the same number as Sec. 2 ) through the Galaxy. The injected and observed skymaps obtained are used to determine the dipole and quadrupole amplitudes given by Eq. 3 for this setup, which are random in origin. We repeated this $10^{4}$ times and fit the distribution of the dipole and the quadrupole with a Gaussian. The 99.5% upper limit on the noise dipole and quadrupole amplitudes from this distribution provide a limit below which any amplitude calculated is considered to be at the noise level. Analytically, this estimate on the noise level for each multipole can be estimated by $3\sqrt{2^{l}/N_{\rm{obs}}}$, where $l$ is the multipole number and $N_{\rm obs}$ is the number of particles in the UHECR skymap.
The observed skymaps for cases 1, 2, and 3 are shown in the top-right, top-left and bottom-left panel of Fig. 4, respectively. Similar to the injected skymaps the NSIDE = 512 and they have been smoothed by a top-hat function with an angle of $45^{\circ}$ (Abreu et al., 2022; Abdul Halim et al., 2024). For the case when the diffusive sphere in which the UHECRs propagate is comparable or larger than the scattering length of the UHECRs, a large-scale dipolar skymap is obtained. Similar to the injected case we remove the monopole for the extragalactic dipole, single and three-beam observed skymaps. In the three beam case (see bottom-left panel in Fig. 4) the choice of our weights given to the three sources results in observed dipole direction not very far from what is observed by PAO (Abdul Halim et al., 2024). The power spectrum for the three cases is shown in the bottom-right panel of Fig. 4.
In Fig. 5 we examine the observed dipole and quadrupole amplitudes for the three distinct source scenarios and show the following:
@@
    Left panel of Fig. 5: The observed dipole amplitude, $\delta_{0}$, is given by the hollow markers in red, green, and blue for the extragalactic dipole, single and three-beam cases. The dipole is seen to become gradually suppressed as the scattering length, $L_{\rm{scat}}$, is decreased. The solid markers denote the observed dipole amplitude, $\delta_{\rm{all}}$, colour-coded as above, in this case. For the single and three beam cases, starting at small $L_{\rm{scat}}$, as $L_{\rm{scat}}$ is increased the dipole amplitude initially rises and then falls. The growth in the dipole amplitude results from a reduction in diffusive ”optical depth” of the halo ($R_{\rm I}/L_{\rm{scat}}$). The subsequent decrease in the dipole ($\delta_{\rm{all}}$) at larger $L_{\rm scat}$ values results from power migrating from smaller to larger $\ell$ as $L_{\rm{scat}}$ is increased, and the corresponding probability that particles in the volume scatter before escaping decreases. In appendix A, we plot the power spectrum for three different scattering length values in Fig. 6 to demonstrate this point. In the extragalactic dipole case, the solid red markers ($\delta_{\rm{all}}$) behave similarly to hollow red markers ($\delta_{0}$), as power at large values of $\ell$ (>> 1) is at the noise level. Knowledge of the dipole amplitude alone is insufficient to distinguish between the different source scenarios considered.
@@
    Right panel of Fig. 5. The observed quadrupole amplitude, $\epsilon_{0}$, is given by the hollow markers and $\epsilon_{\rm{all}}$ is given by the solid markers in red, green, and blue for the extragalactic dipole, single and three-beam cases. For the extragalactic dipole scenario only a dipole is injected, so no significant quadrupole amplitude is expected to be observed. A non-zero quadrupole amplitude is only obtained for the single and three beam cases. For these beam scenarios, starting from small $L_{\rm{scat}}$, as $L_{\rm{scat}}$ is increased the quadrupole amplitude ($\epsilon_{\rm{all}}$) initially rises and then falls. This behaviour occurs for similar reasons as for the dipole behaviour.
@@ 4 Discussion
Figure 5: Left-hand panel: dipole amplitude $\delta^{\rm{obs}}$ as a function of scattering length for the three cases: 1) extragalactic dipole shown in red ($\star$), 2) single-beam case shown in green ($\blacksquare$) and 3) three-beam case shown in blue ( ). The solid and hollow markers indicate $\delta_{\rm{all}}$ and $\delta_{0}$, respectively, for all three cases. Right-hand panel: quadrupole amplitude for the three cases as on the left figure with the same colour code, here $\epsilon_{\rm{all}}$ and $\epsilon_{0}$ are shown by solid and hollow markers again. The orange band represents the $99.5\%$ confidence region for the noise. The vertical dashed line indicates where $L_{\rm{scat}}\approx R_{\rm I}$.
@@ 5 Conclusions
Recent observations of emission from the halo of nearby galaxies suggest that a considerable thermal gas pressure fills galactic haloes out to distances beyond 100 kpc from the galactic centres (Martynenko, 2022; Bregman et al., 2022; Zhang et al., 2024). Here, we consider the consequences of these observations on extragalactic UHECR propagation through the Milky Way’s giant Galactic halo, assuming that the halo also possesses a considerable magnetisation fraction. We focus on the effects of a turbulent magnetic field in the halo. To explore a range of possibilities in this direction, we consider three distinct injection scenarios, namely: $\textit{1})$ an extragalactic dipole, 2) a single beam (a small spread of arrival directions around a single direction) and 3) three beams.
@@ Acknowledgements
VS and AT acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. AvV acknowledges support from Khalifa University’s internal FSU-2022-025 and RIG-2024-047 grants. VS would like to thank Clive Dickinson and the STFC Consolidated Grant (ST/P000649/1) for supporting part of the work carried out at the University of Manchester. The authors would also like to thank James Matthews and Damiano Fiorillo for useful discussions.
@@ Appendix A Power spectrum for different $L_{\rm{scat}}$ values
We plot the power spectrum from the observed skymaps for three different values of $L_{\rm{scat}}$ in Fig. 6 for the three beam case with $L_{\rm{coh}}\approx$ 1 kpc, most of the power sits in the monopole, dipole and quadrupole.